OpenStudy (calculusxy):

MEDAL FOR ANYONE WHO HELPS!!! Last spring, Hana finished a 15 kilometer bicycle race in 3/4 hour. This fall, she wants to ride in an 18.6 mile bicycle race. She set a goal to increase her average speed in miles per hour, for the 18.6 mile by 25% compared to his average speed in the 15 kilometer race. If Hana meets her goal, in what amount of time, in hours and minutes, will she finish the 18.6 mile race?

OpenStudy (calculusxy):

@ganeshie8

OpenStudy (perl):

so her speed last fall was 15 km / (3/4) hr

OpenStudy (calculusxy):

yes and 3/4 hour is equal to 15 minutes right?

OpenStudy (perl):

45

OpenStudy (calculusxy):

i am sorry. i had a typo.

OpenStudy (perl):

no worries :)

OpenStudy (calculusxy):

so how can i solve this problem?

OpenStudy (perl):

so we want to increase her average speed. first I would change the average speed of fall to miles / hr

OpenStudy (perl):

lets keep hours as hours, so 3/4 hours = .75 hour

OpenStudy (calculusxy):

okay.

OpenStudy (perl):

1 mile = 1.60934 km

OpenStudy (perl):

you can derive this 1 inch = 2.54 cm (exactly) 1 inch = .0254 meter 1 foot = 12 * .0254 meters 1 mile = 5280 * 12 * .0254 meters 1 mile = 5280 * 12 * .0254 / 1000 kilometers

OpenStudy (perl):

A mile is exactly 1.609344 kilometers

OpenStudy (calculusxy):

so do i do 18.6 times 1.609344?

OpenStudy (perl):

(15 km / .75 hr ) * ( 1 mi / 1.609344 km )

OpenStudy (calculusxy):

this question is so confusing to me and i have trying to solve this for the past 30 minutes, yet, i have not gotten a perfect solution for this.

OpenStudy (perl):

we will get it, just bear with me.

OpenStudy (perl):

ok change average speed in fall to miles per hour (15 km / .75 hr ) * ( 1 mi / 1.609344 km ) = 12.427 miles per hour

OpenStudy (calculusxy):

wait did you do: \[\frac{ 15km }{ 0.75hr } \times \frac{ 1mi }{ 1.609344km } = \]

OpenStudy (perl):

yes :)

OpenStudy (calculusxy):

i don't know how you got 12.427 mph

OpenStudy (perl):

15/ .75 = 20 , right?

OpenStudy (perl):

then 20 / 1.609344 = 12.427

OpenStudy (calculusxy):

okay and then?

OpenStudy (perl):

i think its easier if we go back and do this differently. lets change the 18.6 miles to km

OpenStudy (calculusxy):

okay?

OpenStudy (perl):